OpenStudy (anonymous):

r= 8 sin theta to rectangular form

OpenStudy (anonymous):

r= 8 sin theta multiply both sides by r we have r^2 = 8rsin theta

OpenStudy (anonymous):

now use r^2 =x^2+y^2 and rcos theta =x and rsin theta =y